# Minchul Kim, Anil Jain, Xiaoming Liu

### v2

@@ 4.3 Neural Network Architectures
CNN Architectures in FR: The backbone neural network architecture plays a crucial role in extracting discriminative features from face images. The revolution brought by deep learning in computer vision, largely initiated by AlexNet [15] on the ImageNet challenge, quickly permeates the field of FR. Early deep FR models adapt existing successful CNN architectures designed for general object recognition.
@@
Architectures like GoogLeNet [95] demonstrates the power of increased network depth and led to its adoption in FaceNet [14]. The introduction of Residual Networks (ResNets) [16] which addresses the vanishing gradient problem in very deep networks through the use of residual connections (shortcuts) leads to training of much deeper models (e.g., ResNet-50, ResNet101, ResNet-152). Variants of ResNet (e.g., SE blocks [96]), become the popular backbone for many SoTA face recognition systems developed in the late 2010s [20, 19]. ArcFace [20]’s adoption of input size $112\times 112$ leads to the widely used IR-ResNet face recognition backbones which removed first downsampling blocks to compensate for the small resolution. Fig. 12 shows the progression of facial image sizes in the FR datasets.
@@
Facial alignment is a crucial preprocessing step in FR systems, ensuring that key facial features (i.e. eyes, nose, and mouth) are consistently positioned across different images. Earlier FR datasets [53] utilize Multi-task Cascaded Convolutional Network (MTCNN)[97], which jointly performs face detection and landmark localization via a series of cascaded networks. With the advent of single-stage detectors like SSD [98], more efficient and accurate methods emerge. Notably, RetinaFace [99] has become a popular solution, offering precise face detection and alignment. When trained on strong datasets such as WiderFace [100] and paired with an improved backbone, RetinaFace is a robust choice for preprocessing large-scale face datasets [55]. Some example alignments are shown in the last row of Fig. 12.
@@
Figure 11: Comparison of architectures used in FR. From left to right: ResNet [16] is an architecture used for classification [15]. IResNet [20] modifies this by removing downsampling and using feature flattening, batch normalization (BN), and fully connected (FC) layers that are helpful for metric learning [101]. Vision Transformer [102] (ViT) replaces convolutions with a patchify operation and transformer blocks; ViT variants further extend this by incorporating auxiliary information such as facial keypoints [99, 39] to improve learning.
@@
Vision Transformers in FR: Mirroring trends in natural language processing and broader computer vision, Vision Transformers (ViTs) [102] have emerged as a powerful alternative to CNNs. ViTs process images by dividing them into patches, and feeding the resulting sequence into a Transformer encoder [103]. The self-attention mechanism within Transformers allows the model to weigh the importance of different image patches globally, potentially capturing long-range dependencies that might be missed by the local receptive fields of CNNs. ViTs have also shown great performance in face recognition domains [104, 105, 39]. And adoption of ViT in face recognition implies adoption of advances around ViT. SwinFace [105] is an application of Swin Transformer [106]. KP-RPE[39] integrates facial landmarks into relative position encodings in ViT, improving robustness to pose variations.
@@
Empirically, compared to ResNets, training ViT on FR models entails more augmentations and requires larger-scale training set [39]. Fig. 11 shows how FR models have changed over time, from ResNets to newer ViT models that use attention and extra information to recognize faces better.
Figure 12: Comparison of face recognition methods in terms of facial contextual area and image resolution. Models have evolved to focus on more tightly cropped facial regions while reducing resized input image size ($256\times 256$ to $112\times 112$), enabling more efficient feature extraction.
@@
While both CNNs, particularly ResNet variants like IR-SE models, and Vision Transformers (ViTs) have demonstrated SoTA performance, there is no single definitively best architecture for all FR tasks. ViTs have shown potential for marginally higher accuracy on some benchmarks, especially when trained on extremely large datasets, leveraging their capability to capture global image dependencies. However, they often necessitate more extensive training data and sophisticated augmentation strategies. ResNets remain highly competitive, often providing a more efficient balance between performance and training/inference cost, particularly with moderate-sized datasets. The optimal choice often depends on the specific application’s constraints, including dataset size, available compute resources, and deployment environment.
@@
Face recognition model deployment depends heavily on computational demands, mainly measured by FLOPs and model size. For example, IResNet50 needs 12.62 GFLOPs and has 43.59M parameters, while IResNet101 uses 24.19 GFLOPs and 65.15M parameters. ViT models are more demanding—ViT Small has 17.42 GFLOPs and 95.95M parameters, and ViT Base requires 24.83 GFLOPs with 114.87M parameters. On consumer GPUs such as Nvidia 3090, IResNet50 can process over 1400 images/second, while ViT Base handles about 640 images/second. Unlike academic research, industry models or government vendor models [107] can use model ensembles, further increasing the load. Additionally, pre-processing steps like face detection and alignment add to the overall computational cost. Note that ViT backbones can benefit from research that speed up ViT inference [108, 109].
Efficiency, Adaptation, and Compact Embeddings: Beyond achieving maximum accuracy, research has also focused on developing efficient architectures suitable for deployment on resource-constrained devices like mobile phones. MobileFaceNet [110] employs depthwise separable convolutions to significantly reduce computational cost and model size while maintaining reasonable accuracy (IResNet100 at 99.83% vs MobileFaceNet at 99.55% in LFW verification accuracy while being 60$\times$ smaller). S-ViT[111] applies sparse attention to reduce computational cost without sacrificing accuracy. The continuous evolution of neural network architectures, from deeper CNNs to attention-based Transformers and efficient mobile designs, has been a key driver alongside loss function innovations and larger datasets in pushing the performance boundaries of automated FR.
@@
Fine-tuning is crucial for adapting face recognition models to new domains, especially under quality mismatches [107] (e.g. low vs high quality images). Instead of full fine-tuning, which risks catastrophic forgetting, recent work like PETALface [112] uses LoRA [113], a parameter-efficient finetuning method that adds low-rank adaptation modules. By weighting LoRA blocks based on image quality, PETALface adapts effectively to low-resolution faces while preserving high-resolution performance.
@@
TABLE III: Comparison of synthetic face training datasets for face recognition across five standard benchmarks. ”Gap to Real” shows the average performance drop relative to the use of real CASIA-WebFace dataset alone for training. Brackets in the Generator Training Dataset column denote datasets used for pretraining, which may help models learn facial priors. A more fair comparison might involve equalizing the use of pretrained models. FR model used is IR50 [114, 115, 116].
Methods Venue Generator Train Dataset # images (# IDs×$\times$ imgs/ID) LFW CFP-FP CPLFW AgeDB CALFW Avg Gap to Real
SynFace [114] ICCV21 FFHQ [117] 0.5M (10K $\times$ 50) 91.93 75.03 70.43 61.63 74.73 74.75 26.04
DigiFace [115] WACV23 511 3D Scans 1M (10K $\times$ 100) 95.40 87.40 78.87 76.97 78.62 83.45 11.34
DCFace [116] CVPR23 CASIA-WebFace (FFHQ) 0.5M (10K $\times$ 50) 98.55 85.33 82.62 89.70 91.60 89.56 5.23
IDnet [118] CVPR23 CASIA-WebFace [53] 0.5M (10K $\times$ 50) 92.58 75.40 74.25 63.88 79.90 79.13 15.66
ExFaceGAN [119] IJCB23 CASIA-WebFace 0.5M (10K $\times$ 50) 93.50 73.84 71.60 78.92 82.98 80.17 14.62
SFace2 [120] TBIS24 CASIA-WebFace 0.6M (10K $\times$ 60) 96.50 77.11 74.60 77.37 83.40 81.62 13.17
Arc2Face [121] ECCV24 WF42M [55] (Stable Diffusion) 0.5M (10K $\times$ 50) 98.81 91.87 85.16 90.18 92.63 91.73 3.06
Vec2Face [122] ICLR2025 CASIA-WebFace (WebFace4M) 0.5M (10K $\times$ 50) 98.87 88.97 85.47 93.12 93.57 92.00 2.79
Alongside the choice of network architecture, the dimensionality of the output feature vector, or template size, is another critical design decision influencing system performance and efficiency. While modern systems often employ 512-dimensional embeddings, research by Gong et al. [123] investigates the fundamental requirements of these representations by examining their intrinsic dimensionality. Their work reveals that the actual degrees of freedom needed to capture essential facial variations (the intrinsic dimension) can be substantially lower than the commonly used ambient dimensions (e.g., d=512). Also FaceNet [14] indicates that for a given training dataset, higher performance was achieved with d=128 compared to d=512. This finding implies that standard high-dimensional face embeddings contain significant redundancy, suggesting the potential to develop much more compact templates that could enable faster search and more efficient storage while retaining discriminative power. It is important to note, however, that in commercial systems, a template often includes more than just the feature vector itself. It may contain metadata such as quality indicators, detection attributes, or even multiple embeddings, leading to greater variation in template sizes.
4.4 Synthetic Datasets
The growing demand for large-scale, diverse, and ethically sourced training datasets has driven increasing interest in the use of synthetic face data. Collecting real-world facial datasets often introduces privacy and consent challenges, as well as issues of demographic imbalance and limited representation under challenging conditions (e.g., extreme poses, rare ethnicities). Synthetic data offers a compelling alternative or data augmentation by enabling controlled, scalable, and bias-aware dataset creation [114, 116, 115].
The generation of photorealistic synthetic faces has been significantly advanced by deep generative models, particularly Generative Adversarial Networks (GANs)[124]. Variants such as StyleGAN[117, 125] are especially effective at producing high-resolution facial imagery, capable of modeling complex visual distributions and allowing fine-grained control over attributes like pose, expression, and illumination via manipulations in the latent space. The integration of 3D Morphable Models (3DMM) into GANs has further enhanced the controllability of facial attributes during generation [126, 127, 128, 129]. For instance, CFSM [130] leverages GANs to synthesize faces with diverse styles, aiding in the generation of richly varied datasets.
@@ 4.4 Synthetic Datasets
Recently, Diffusion Models [131] emerge as a powerful generative paradigm, achieving impressive image quality and diversity. They generate images by gradually denoising a sample from pure noise, learning to reverse the diffusion process. Text-conditioned diffusion models are especially effective for controlled synthesis, enabling detailed and semantically guided generation [132, 133]. ControlNet [134] and IP-Adapter [135] make the model adhere to input conditions such as facial landmarks, masks or other clues.
@@
Leveraging these generative capabilities, researchers have explored creating entire synthetic datasets specifically for training FR models. The goal of face dataset generation is to create multiple images of the same subject at a large scale. The ID consistency is at an interplay with the diversity of generated images. SynFace [114] first applies GAN and latent interpolation method to generate face datasets, resulting in average face verification rate of $74.75\%$, marking a significant drop compared to real CASIA-Webface dataset of $94.79\%$ as in Tab. III. Since then, mulitple works have attempted to reduce the gap.
@@
Following SynFace, the field has witnessed rapid growth, with numerous efforts aiming to bridge the performance gap between synthetic and real face datasets. One notable direction is the use of diffusion-based models, as exemplified by DCFace [116], which separates identity and style conditions during generation to produce identity-consistent subjects while maintaining high diversity. Arc2Face [121] takes this further by leveraging a pretrained Stable Diffusion model, highlighting the power of foundation models in capturing generalizable facial representations. On the other hand, Vec2Face [122] demonstrates that GAN-based synthesis can remain competitive when guided by a FR feature space, emphasizing the critical role of identity disentanglement. These approaches reflect the ongoing trend in synthetic dataset generation: pushing the frontier of realism and diversity while preserving identity consistency to close the gap with real-world data. Tab. III shows the performance of recent synthetic face datasets on FR.
There is growing interest in synthetic FR challenges, as reflected by the FRCSyn (Face Recognition Challenge in the Era of Synthetic Data) series [136, 137]. The series reveal that while synthetic-only training remains slightly behind real data in certain scenarios, it offers compelling advantages in mitigating demographic bias and improving robustness to variations in pose, age, and occlusion. Examples of the comparison between real and synthetic datsets are shown in Fig. 13. Also multiple studies show that snynthetic images can be used as an augmentation to the real dataset. For example, addition of Vec2Face [122] to CASIA-WebFace can increase about 1.00% in average verification accuracy.
Despite the promise, the effective use of purely synthetic data for training SoTA FR models faces challenges. A primary hurdle is the domain gap between synthetic and real images. Models trained solely on synthetic data may not generalize very well to real-world images due to subtle differences in texture, lighting, or artifacts introduced by the generation process. Ensuring sufficient diversity and realism, particularly in capturing the nuances of identity across various conditions, remains an active research area. Furthermore, the generation process itself can be computationally expensive, and ethical concerns regarding the potential misuse of highly realistic synthetic faces (e.g., deepfakes) persist. To address these risks, there is a growing emphasis on watermarking or other techniques [138] to clearly indicate that images are synthetically generated.

### v3

@@ 4.3 Neural Network Architectures
CNN Architectures in FR: The backbone neural network architecture plays a crucial role in extracting discriminative features from face images. The revolution brought by deep learning in computer vision, largely initiated by AlexNet [krizhevsky2012imagenet] on the ImageNet challenge, quickly permeates the field of FR. Early deep FR models adapt existing CNN architectures designed for general object recognition.
@@
Architectures like GoogLeNet [szegedy2015going] demonstrate the power of increased network depth and led to its adoption in FaceNet [schroff2015facenet]. The introduction of Residual Networks (ResNets) [he2016deep] which addresses the vanishing gradient problem in very deep networks through the use of residual connections (shortcuts) leads to training of much deeper models (e.g., ResNet-50, ResNet101, ResNet-152). Variants of ResNet (e.g., SE blocks [hu2018squeeze]), become the popular backbone for many SoTA FR systems developed in the late 2010s [deng2019arcface, wang2018cosface]. ArcFace [deng2019arcface]’s adoption of input size $112\times 112$ leads to the widely used IR-ResNet backbones which removed first downsampling blocks to compensate for the small resolution. Fig. 12 shows the progression of facial image sizes in the FR datasets.
@@
Facial alignment is a crucial preprocessing step in FR systems, ensuring that key facial features (i.e. eyes, nose, and mouth) are consistently positioned across different images. Earlier FR datasets [casia] utilize Multi-task Cascaded Convolutional Network (MTCNN)[zhang2016joint], which jointly performs face detection and landmark localization via a series of cascaded networks. With the advent of single-stage detectors like SSD [liu2016ssd], more efficient and accurate methods emerge. Notably, RetinaFace [deng2019retinaface] has become a popular solution, offering precise face detection and alignment. When trained on strong datasets such as WiderFace [yang2016wider] and paired with an improved backbone, RetinaFace is a robust choice for preprocessing large-scale face datasets [webface]. Some example alignments are shown in the last row of Fig. 12.
@@
Figure 11: Comparison of architectures used in FR. From left to right: ResNet [he2016deep] is an architecture used for classification [krizhevsky2012imagenet]. IResNet [deng2019arcface] modifies this by removing downsampling and using feature flattening, batch normalization (BN), and fully connected (FC) layers that are helpful for metric learning [luo2019bag]. Vision Transformer [dosovitskiy2020image] (ViT) replaces convolutions with a patchify operation and transformer blocks; ViT variants further extend this by incorporating auxiliary information such as facial keypoints [deng2019retinaface, kim2024keypoint] to improve learning.
@@
Vision Transformers in FR: Mirroring trends in natural language processing and broader computer vision, Vision Transformers (ViTs) [dosovitskiy2020image] have emerged as a powerful alternative to CNNs. ViTs process images by dividing them into patches, and feeding the resulting sequence into a Transformer encoder [vaswani2017attention]. The self-attention mechanism within Transformers allows the model to weigh the importance of different image patches globally, potentially capturing long-range dependencies that might be missed by the local receptive fields of CNNs. ViTs have also shown great performance in FR domains [rodrigo2024comprehensive, qin2023swinface, kim2024keypoint]. And adoption of ViT in FR implies adoption of advances around ViT. SwinFace [qin2023swinface] is an application of Swin Transformer [liu2021swin]. KP-RPE[kim2024keypoint] integrates facial landmarks into relative position encodings in ViT, improving robustness to pose variations.
@@
Empirically, compared to ResNets, training ViT on FR models entails more augmentations and requires larger-scale training set [kim2024keypoint]. Fig. 11 shows how FR models have changed over time, from ResNets to newer ViT models that use attention and extra information to recognize faces better.
@@
While both CNNs, particularly ResNet variants like IR-SE models, and Vision Transformers (ViTs) have demonstrated SoTA performance, there is no single definitively best architecture for all FR tasks. ViTs have shown potential for marginally higher accuracy on some benchmarks, especially when trained on extremely large datasets, leveraging their capability to capture global image dependencies. However, they often necessitate more extensive training data and sophisticated augmentation strategies. ResNets remain highly competitive, often providing a more efficient balance between performance and training/inference cost, particularly with moderate-sized datasets. The optimal choice often depends on the specific application’s constraints, including dataset size, compute resources, and deployment setting.
@@
FR model deployment depends heavily on computational demands, mainly measured by FLOPs and model size. For example, IResNet50 needs 12.62 GFLOPs and has 43.59M parameters, while IResNet101 uses 24.19 GFLOPs and 65.15M parameters. ViT models are more demanding—ViT Small has 17.42 GFLOPs and 95.95M parameters, and ViT Base requires 24.83 GFLOPs with 114.87M parameters. On consumer GPUs such as Nvidia 3090, IResNet50 can process over 1400 images/second, while ViT Base handles about 640 images/second. Unlike academic research, industry models or government vendor models [liu2024farsight] can use model ensembles, further increasing the load. Preprocessing steps like face detection and alignment add to the computational cost. Note that ViT backbones can benefit from research that speed up ViT inference [dao2022flashattention, xFormers2022].
Figure 12: Comparison of FR in terms of facial contextual area and image resolution. Models have evolved to focus on more tightly cropped regions while reducing resized input image size ($256\times 256$ to $112\times 112$), enabling more efficient feature extraction.
[…]
@@
Fine-tuning is crucial for adapting FR models to new domains, especially under quality mismatches [liu2024farsight] (e.g. low vs high quality images). Instead of full fine-tuning, which risks catastrophic forgetting, recent work like PETALface [narayan2025petalface] uses LoRA [hu2022lora], a parameter-efficient finetuning method that adds low-rank adaptation modules. By weighting LoRA blocks based on image quality, PETALface adapts effectively to low-resolution faces while preserving high-resolution performance.
TABLE II: Comparison of synthetic face training datasets for FR across five standard benchmarks. ”Gap to Real” shows the average performance drop relative to the use of real CASIA-WebFace dataset alone for training. Brackets in the Generator Training Dataset column denote datasets used for pretraining, which may help models learn facial priors. A more fair comparison might involve equalizing the use of pretrained models. FR model used is IR50 [qiu2021synface, bae2023digiface, kim2023dcface].
[…]
SynFace [qiu2021synface] ICCV21 FFHQ [karras2019style] 0.5M (10K $\times$ 50) 91.93 75.03 70.43 61.63 74.73 74.75 26.04
DigiFace [bae2023digiface] WACV23 511 3D Scans 1M (10K $\times$ 100) 95.40 87.40 78.87 76.97 78.62 83.45 11.34
DCFace [kim2023dcface] CVPR23 CASIA-WebFace (FFHQ) 0.5M (10K $\times$ 50) 98.55 85.33 82.62 89.70 91.60 89.56 5.23
IDnet [kolf2023identity] CVPR23 CASIA-WebFace [casia] 0.5M (10K $\times$ 50) 92.58 75.40 74.25 63.88 79.90 79.13 15.66
ExFaceGAN [boutros2023exfacegan] IJCB23 CASIA-WebFace 0.5M (10K $\times$ 50) 93.50 73.84 71.60 78.92 82.98 80.17 14.62
SFace2 [boutros2024sface2] TBIS24 CASIA-WebFace 0.6M (10K $\times$ 60) 96.50 77.11 74.60 77.37 83.40 81.62 13.17
Arc2Face [papantoniou2024arc2face] ECCV24 WF42M [webface] (Stable Diffusion) 0.5M (10K $\times$ 50) 98.81 91.87 85.16 90.18 92.63 91.73 3.06
Vec2Face [wu2024vec2face] ICLR2025 CASIA-WebFace (WebFace4M) 0.5M (10K $\times$ 50) 98.87 88.97 85.47 93.12 93.57 92.00 2.79
Also, FaceNet [schroff2015facenet] indicates that for a given training dataset, higher performance was achieved with d=128 compared to d=512. This finding implies that standard high-dimensional face embeddings contain significant redundancy, suggesting the potential to develop much more compact templates that could enable faster search and more efficient storage while retaining discriminative power.
The optimal embedding dimensionality varies with the scale and diversity of the training data. The embedding space must fit all subjects within a hypersphere while preserving sufficient inter-class distances. As the number of identities increases, a higher-dimensional feature space may be needed to maintain discriminability. Hence, smaller dimensions (e.g., 128) may suffice for moderate datasets, whereas larger datasets benefit from higher-dimensional representations to preserve subject separation.
4.4 Synthetic Datasets
The growing demand for large-scale, diverse, and ethically sourced training datasets has driven increasing interest in the use of synthetic face data. Collecting real-world facial datasets often introduces privacy and consent challenges, as well as issues of demographic imbalance and limited representation under challenging conditions (e.g., extreme poses, rare ethnicities). Synthetic data offers a compelling alternative or data augmentation by enabling controlled, scalable, and bias-aware dataset creation [qiu2021synface, kim2023dcface, bae2023digiface].
The generation of photorealistic synthetic faces has been significantly advanced by deep generative models, particularly Generative Adversarial Networks (GANs)[goodfellow2014generative]. Variants such as StyleGAN[karras2019style, Karras2019stylegan2] are especially effective at producing high-resolution facial imagery, capable of modeling complex visual distributions and allowing fine-grained control over attributes like pose, expression, and illumination via manipulations in the latent space. The integration of 3D Morphable Models (3DMM) into GANs has further enhanced the controllability of facial attributes during generation [deng2018uv, gecer2018semi, geng20193d, kim2018deep]. For instance, CFSM [liu2022controllable] leverages GANs to synthesize faces with diverse styles, aiding in the generation of richly varied datasets.
@@ 4.4 Synthetic Datasets
Recently, Diffusion Models [ho2020denoising] emerge as a powerful generative paradigm, achieving impressive image quality and diversity. They generate images by gradually denoising a sample from pure noise, learning to reverse the diffusion process. Text-conditioned diffusion models are especially effective for controlled synthesis, enabling detailed and semantically guided generation [dhariwal2021diffusion, rombach2022high]. ControlNet [controlnet] and IP-Adapter [ipadapter] make the model adhere to input conditions such as facial landmarks, masks or other clues.
@@
Leveraging these generative capabilities, researchers have explored creating entire synthetic datasets specifically for training FR models. The goal of face dataset generation is to create multiple images of the same subject at a large scale. The ID consistency is at an interplay with the diversity of generated images. SynFace [qiu2021synface] first applies GAN and latent interpolation method to generate face datasets, resulting in average face verification rate of $74.75\%$, marking a significant drop compared to real CASIA-Webface dataset of $94.79\%$ as in Tab. II. Since then, mulitple works have attempted to reduce the gap.
@@
Following SynFace, rapid advances have sought to bridge the gap between synthetic and real face datasets. Diffusion-based models such as DCFace [kim2023dcface] separate identity and style conditions to produce identity-consistent and diverse subjects, while Arc2Face [papantoniou2024arc2face] builds on pretrained Stable Diffusion to exploit the generalization power of foundation models. Vec2Face [wu2024vec2face] shows that GAN-based synthesis remains competitive when guided by a FR feature space, underscoring the importance of identity disentanglement. These methods exemplify the trend toward improving realism, diversity, and identity consistency to narrow the gap with real data. Tab. II presents the performance of recent synthetic face datasets on FR.
There is growing interest in synthetic FR challenges. FRCSyn series [melzi2024frcsyn, deandres2024frcsyn], show that while synthetic-only training trails real data slightly, it reduces demographic bias and improves robustness to pose, age, and occlusion. Figure 13 compares real and synthetic datasets. Adding synthetic data, such as Vec2Face [wu2024vec2face] to CASIA-WebFace, can raise average verification accuracy by about 1.00%.
Despite its promise, using purely synthetic data to train SoTA FR models faces key challenges, primarily the domain gap between synthetic and real images. Models trained only on synthetic data often struggle to generalize to real-world due to subtle differences in texture, lighting, or artifacts from the generation process. Achieving sufficient diversity and realism, especially in capturing identity nuances under varying conditions, remains an active research focus. Generative models can mitigate privacy and consent concerns, yet their training on real web-sourced images raises doubts about whether generated content removes or only obscures data ownership and consent issues [Silberling2025]. To mitigate these risks, growing attention is given to watermarking and related methods [asnani2022proactive] that clearly mark images as synthetic.